OpenStudy (anonymous):

How do i determine the value? 1. sec 15 degree 30' 2. cot 74 degree 10' 3. sin 75 degree 50'

OpenStudy (anonymous):

\[\text{Sec}\left[\left(15+\frac{30}{60}\right){}^{\circ}\right]=1.03774 \]\[\text{Cot}\left[\left(74+\frac{10}{60}\right){}^{\circ}\right]=0.2836 \]\[\text{Sin}\left[\left(75 +\frac{50}{60}\right){}^{\circ}\right]=0.969588 \]